OpenStudy (anonymous):

log(x2-2x+1)= log(x2+4x-5) i need solution

OpenStudy (anonymous):

None. I think that's what it is. No solutions exist.

OpenStudy (anonymous):

but this is the question

OpenStudy (anonymous):

x^2-2x+1=x^2+4x-5 ==> 6x=6 ==> x=1

OpenStudy (anonymous):

oh yeah that's right, the log would cancel. Right?

OpenStudy (anonymous):

exactly, you may cancel

OpenStudy (anonymous):

too much calculus tonight....lol

OpenStudy (anonymous):

thnxx alot @raheen

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

any time @jinausa

OpenStudy (anonymous):

sorry for my horrible response haha.i know better too thats the sad part

OpenStudy (anonymous):

thtss okkk@lawnphysics

OpenStudy (anonymous):

\[\log(x^2-2x+1)= \log(x^2+4x-5)\]\[x^2-2x+1=x^2+4x-5\]\[6x=6\]\[x=1\]